OpenStudy (babyslapmafro):

Could someone please point-out the mistake(s) I am making while trying to find the norm of the following vectors. u=i-3j+2k v=i+j w=2i+2j-4k I am asked to find: ||3u-5v+w|| (click to see my work)

OpenStudy (anonymous):

Are you finding the magnitude?

OpenStudy (babyslapmafro):

=||3(i-3j+2k)-5(i+j)+(2i+2j-4k)|| =||(3i-9j+6k)-(5i+5j)+(2i+2j-4k)|| =||-2j+2k|| =2sqrt(2)

OpenStudy (babyslapmafro):

Answer in book: 2sqrt(37)

OpenStudy (anonymous):

Are you finding the magnitude or not?

OpenStudy (babyslapmafro):

Yes I am asked to find the norm, as stated in the original post. Norm=magnitude=length.

OpenStudy (jack1):

absolute value of 2k - 12j...?

OpenStudy (anonymous):

Notice that:\[\bf 3u-5v+w=[3,-9,6]+[-5,-5,0]+[2,2,-4]=[0,-12,2]\]Then the magnitude is:\[\bf \sqrt{(0)^2+(-12)^2+(2)^2)}=\sqrt{148}=\sqrt {4*37}=2\sqrt{37}\] @Babyslapmafro

OpenStudy (babyslapmafro):

Ok, accidentally added the 5 of the second vector instead of subtracting, thanks...